OpenStudy (nickymail48):

write equation of the line with slope write final answer y=mx+b form: m=-4;(2,-3)

OpenStudy (anonymous):

M=-4 now u need to find the y intercept b

OpenStudy (anonymous):

To solve this, you will need to figure out what B, or the y-intercept, is. This can be done in two ways. The easier way is by putting the equation into point-slope form (y-y1=m(x-x1)) and solving that until you reach slope-intercept form (y=mx+b)

OpenStudy (nickymail48):

so y=-8x+12? or -3x+2y=-4

OpenStudy (nickymail48):

y=-4x+2?

OpenStudy (anonymous):

It would be y=-4x+5 Plug it into point-slope form, then solve that until you reach slope-intercept form. y-y1=m(x-x1) y+3=-4(x-2) y+3=-4x+8 y=-4x+5

OpenStudy (nickymail48):

thanks